# Domain Consistency Representation Learning for Lifelong Person Re-Identification

### v2

@@ II-E Attribute-oriented Anti-Forgetting
II-F Knowledge Consolidation
Maximizing intra-domain discrimination and minimizing inter-domain gaps are in a contradictory relationship. Therefore, achieving a balance between them is crucial for improving the performance of LReID models. Thus,we propose a knowledge consolidation (KC) strategy that leverages global representations for knowledge transfer between the old and new models. This includes alignment loss and logit-level distillation loss.
Maintaining distribution consistency between the old and new models for previous datasets can limit the model’s ability to learn new knowledge. Therefore, we propose an alignment loss to explore global representations for knowledge transfer from the current dataset, as follows:
We further introduce a logit-level distillation loss to enhance the extraction of identity information shared between the old and new models, further improving the model’s knowledge consolidation ability. This is represented as follows:
The knowledge consolidation loss is defined as:
The total loss function is formulated as:
III Experiments
III-A Experiments Setting
@@
Datasets. To verify the performance of our method in anti-forgetting and generalization, we evaluate our method on a challenging benchmark consisting of Market1501 [38], CUHK-SYSU [39], DukeMTMC [40], MSMT17V2 [41] and CUHK03 [42], referred to as seen datasets. Two representative training orders are set up following the protocol described in [17] for training and testing. Further, we employ six datasets including VIPeR [43], GRID [44], CUHK02 [45], OccDuke [46], OccREID [47], and PRID2011 [48], as unseen dataset.
@@
Implementation Details. Our text encoder and image encoder are based on a pre-trained CLIP model, while the attribute decoder utilizes a transformer-based architecture[36]. All person images are resized to 256128. We use Adam [49] for optimization and train each task for 60 epochs. The batch size is set to 128. The learning rate is initialized at 5 and is decreased by a factor of 0.1 every 20 epochs for each task. We employ mean average precision (mAP) and Rank-1 accuracy (R-1) to evaluate the LReID model on each dataset.
III-B Comparison with SOTA Methods
@@
We compare the proposed DCR with SOTA LReID to demonstrate the superiority of our method, including AKA[17], PTKP[1], PatchKD[14], KRKC[2], and ConRFL[9], DKP[15], C2R[16], LSTKC[13]. Experimental results on training order-1 and order-2 are shown in Table I and Table II, respectively. CODA [50] method employs ViT-B/16 as the backbone.
@@
Compared with LReID methods. In Table I and Table II, Our DCR significantly outperforms LReID methods, with an seen-avg incremental gain of 10.0% mAP/7.8% R-1, and 9.8% mAP/7.5% R-1 on training order-1 and order-2, respectively. Meanwhile, our DCR effectively alleviate catastrophic forgetting, achieving 6.9% mAp/1.1% R-1, and 5.4% mAP/ 2.2% R-1 improvement on the first dataset (Mrket1501 and DukeMTMC) with different training orders. Compared to CODA, our DCR significantly outpreform performance under the backbone of VIT-B/16. Additionally, our DCR improves the average by 8.1 mAP%/7.5% R-1 and 9.5% mAP/11.0% R-1 on unseen datasets. In contrast, our DCR achieves a trade-off between anti-forgetting and acquiring new information, significantly enhances generalization capabilities.
@@
Compared with Baseline. Due to the lack of CLIP-based comparison methods in LReID, we introduce a Baseline model including CLIP model, attribute-text generator and knowledge consolidation strategy. Compared to the Baseline, Our DCR improves the Seen-Avg by 11.4% mAP/10.4% R-1 and by 9.8% mAP/10.2% R-1. These results demonstrate that our proposed attribute-wise representations learning achieves significant performance in balancing maximization intra-domain discrimination and minimization inter-domain gaps in LReID.
The effectiveness of minimizing inter-domain gaps. We visualize the feature distribution of PTKP, KRKC, DKP, and our method across five datasets, as shown in Figure 3. DKP shows poor performance in bridging inter-domain gaps, as knowledge prototypes struggle to fit the data distribution. Compared to other methods, our DCR model better narrows the distribution across increasing datasets. Thus, the proposed DCR model can effectively bridge domain gaps, improving knowledge transfer capabilities, benefiting from the attribute-oriented anti-forgetting (AF) strategy based on attribute-wise representations.
@@
The effectiveness of maximizing intra-domain discrimination. We visualize the feature distribution of KRKC and our method. Figure 4 shows that our DCR model can significantly cluster images of the same identity more tightly (circle) and increase the distance between different identities (black bidirectional arrow). Compared to KRKC, our DCR model effectively improves intra-domain discrimination,
benefiting from the complementary relationship between global and attribute representations, which allows it to learn the subtle nuances of individuals.
@@
Generalization Curves on Unseen Dataset. We analyze the average performance on unseen datasets over the training steps, as shown in Figure 5. Compared to other methods, our DCR model achieves superior performance and exhibits faster performance growth across the training steps. Thus, our attribute-oriented anti-forgetting (AF) strategy effectively bridges inter-domain gaps, enhancing the generalization ability of our model. In summary, our DCR model explores global and attribute-wise representations to achieve a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps.
@@
III-C Ablation Studies
The number of global and attribute-wise repreentations. Global and attribute-wise representations capture individual nuances in intra-domain and inter-domain consistency. We study the suitability of multiple global and attribute-wise representations, as shown in Table III. We observe that setting the number of global and attribute-wise representations to 3 achieves the best performance for our method.
| PFM | ACN | AF | KC | SeenAvg | UnseenAvg | ||
| mAP | R-1 | mAP | R-1 | ||||
| 50.4 | 61.5 | 51.8 | 49.4 | ||||
| 51.7 | 62.1 | 52.5 | 50.3 | ||||
| 57.6 | 68.9 | 58.2 | 56.2 | ||||
| 58.7 | 69.2 | 58.5 | 56.8 | ||||
| 61.8 | 71.9 | 60.8 | 58.3 | ||||
Performance of Different Components. To assess the contribution of each component to our DCR, we conduct ablation studies on seen and unseen datasets, as shown in Table IV. Comparing the first and second rows, we observe that the parallel fusion module (PFM), which employs a parallel cross-attention mechanism, effectively fuses text and image embeddings. Comparing the second and fourth rows, we consider that the attribute compensation network (ACN) and attribute-oriented anti-forgetting (AF) strategy effectively learn domain consistency, improving generalization ability. In the second and third rows, we notice the performance decrease when using only the knowledge consolidation (KC) strategy based on global representations across increasing data while ignoring inter-domain gaps. The results demonstrate that both global representations and attribute-wise representations achieve a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps for enhancing the anti-forgetting and generalization capacity of our DCR.
@@ III-C Ablation Studies
Performance of attribute-text generator. To better understand whether each instance’s text descriptions generated by the attribute-text generator (ATG) provide more fine-grained guidance for learning global representations, we train our model using the generic text descriptor ”A photo of a person” (w/o ATG) for comparison. Table V shows that the attribute-text generator obtain text descriptions to significantly improves overall performance. When using the specific text descriptors, the average decreases by 1.7% mAP/1.4% R-1 on seen datasets and by 1.5% mAP/1.8% R-1 on unseen datasets. ATG enhances the robustness of global representations for each instance, effectively mitigating the forgetting of old knowledge.
IV Conclusion
@@
In this paper, we propose a domain consistency representation learning (DCR) model that explores global and attribute-wise representations to capture subtle nuances in intra-domain and inter-domain consistency, achieving a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps. Specifically, global and attribute-wise representations serve as complementary information to distinguish similar identities in intra-domain. We further develop an attribute-oriented anti-forgetting (AF) strategy and a knowledge consolidation (KC) strategy to minimize inter-domain gaps and facilitate knowledge transfer, enhancing generalization capabilities. Extensive experiments demonstrate that our method achieves superior performance compared to state-of-the-art LReID methods.

### v4

@@ III-F Knowledge Consolidation
IV Experiments
IV-A Experiments Setting
IV-A1 Datasets
@@
To assess the performance of our method in anti-forgetting and generalization, we evaluate our method on a challenging benchmark consisting of Market1501 [42], CUHK-SYSU [43], DukeMTMC [44], MSMT17V2 [45] and CUHK03 [46], referred to as the seen domain. Two representative training orders are set up following the protocol described in [22] for training and testing. Further, we employ six datasets including VIPeR [47], GRID [48], CUHK02 [49], OccDuke [50], OccREID [51], and PRID2011 [52], as the unseen domain. During evaluation, Unseen domain and test sets of the seen domain are combined into a single benchmark. Detailed statistics for these datasets can be shown in TABLE II.
IV-A2 Implementation Details
@@
Our text encoder and image encoder are based on a pre-trained CLIP model, while the attribute decoder utilizes a transformer-based architecture[40]. All person images are resized to 256128. We use Adam [54] for optimization and train each task for 60 epochs. The batch size is set to 128. The learning rate is initialized at 5 and is decreased by a factor of 0.1 every 20 epochs for each task. We employ mean average precision (mAP) and Rank-1 accuracy (R-1) to evaluate the LReID model on each dataset.
IV-B Comparison with SOTA Methods
@@
We compare the proposed DCR with SOTA LReID to demonstrate the superiority of our method, including AKA[22], PTKP[9], PatchKD[16], KRKC[10], ConRFL[12], CODA [53], LSTKC[15], C2R[18], DKP[17]. Experimental results on training order-1 and order-2 are shown in TABLE III and TABLE IV, respectively.
IV-B1 Compared with LReID methods on the Seen Domain
@@
In TABLE III and TABLE IV, our DCR significantly outperforms LReID methods, with an seen-avg incremental gain of 10.0% mAP/7.8% R-1, and 9.8% mAP/7.5% R-1 on training order-1 and order-2, respectively. Meanwhile, our DCR effectively alleviates catastrophic forgetting, achieving 6.9% mAp/1.1% R-1 and 5.4% mAP/2.2% R-1 improvement on the first dataset (Mrket1501 and DukeMTMC) with different training orders. Compared to CODA, our DCR significantly outperforms performance under the backbone of VIT-B/16. In contrast, our DCR achieves a trade-off between anti-forgetting and acquiring new information.
IV-B2 Compared with LReID methods on the Unseen Domain
In Table III, compared to the KRKC method, our DCR model achieves significant improvements, with an average gain of 8.1% mAP / 7.5% R-1 on the unseen domain. Furthermore, when evaluated against the LSTKC methods, our DCR demonstrates superior performance over LReID methods, delivering an unseen-avg incremental improvement of 9.5% mAP/11.0% R-1, as shown in Table IV. We attribute these advancements to the complementary relationship between global and attribute-wise representations, which effectively enhances discrimination among similar identities. Additionally, our attribute-oriented anti-forgetting (AF) and knowledge consolidation (KC) strategies foster inter-domain consistency and facilitate seamless knowledge transfer. In summary, our DCR markedly strengthens generalization capabilities, outperforming existing approaches in terms of mAP and R-1 metrics.
IV-B3 Compared with Baseline
@@
Due to the lack of CLIP-based comparison methods in LReID, we introduce a Baseline model that includes the CLIP model, an attribute-text generator, and a knowledge consolidation strategy. The Baseline outperforms other methods (such as AKA, PTKP, PatchKD, etc.) in mAP and R-1, benefiting from the powerful extraction capabilities of CLIP, as presented in TABLE III and TABLE IV. Compared to the Baseline, our DCR improves the Seen-Avg by 11.4% mAP/10.4% R-1 and by 9.8% mAP/10.2% R-1. These results demonstrate that our proposed domain consistency representation learning strategy achieves significant performance in balancing the maximization of intra-domain discrimination and the minimization of inter-domain gaps in LReID.
IV-B4 The Anti-forgetting Performance of Our Method
We conduct a forgetting measurement experiment in training order-1, as shown in Fig. 3. The Fig. 3 shows the metric measurements for the Market1501 dataset at different training steps. After training on the large-scale MSTMS17 dataset at training step 4, KRKC, LSTCC, and DKP exhibit significant attenuation in mAP and R-1. Because the comparison method limits the performance of the model in minimizing inter-domain gaps. Our method demonstrates a smoother decrease in mAP and Rank-1, which can effectively reduce inter-domain gaps to alleviate the catastrophic forgetting problem.
IV-B5 The effectiveness of minimizing inter-domain gaps
We visualize the feature distribution of PTKP, KRKC, Baseline, and our method across five datasets as shown in Fig. 4. The Baseline shows poor performance in bridging inter-domain gaps, as the lack of attribute-wise representations makes it challenging to reduce inter-domain gaps. The KRKC method effectively separates each domain, but it insufficiently distinguishes identity information within the domain, limiting the model’s ability to prevent forgetting and enhance generalization. Compared to other methods, our DCR effectively distinguishes identity information within a domain and spreads identity information across multiple domains, which significantly improves the anti-forgetting and generalization ability of the model.
IV-B6 The effectiveness of maximizing intra-domain discrimination
@@
We visualize the feature distribution of KRKC and our method. Fig. 5 shows that our DCR can significantly cluster images of the same identity more tightly (circle) and increase the distance between different identities (black bidirectional arrow). Compared to KRKC, our DCR improves intra-domain discrimination due to the complementary relationship between global and attribute-wise representations, which enables it to learn the subtle nuances of individuals.
IV-B7 Generalization Curves on the Unseen Domain
@@
We analyze the average performance on the unseen domain during the training steps, as depicted in Fig. 6. Compared to other methods, our DCR achieves superior performance and exhibits faster performance growth across the training steps. Thus, our attribute-oriented anti-forgetting (AF) strategy effectively bridges inter-domain gaps and enhances the generalization ability of our model. In summary, our DCR explores global and attribute-wise representations to achieve a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps.
@@
| Type | Components | SeenAvg | UnseenAvg | |||||
| PFM | ACN | AF | KC | mAP | R-1 | mAP | R-1 | |
| (a) | 50.4 | 61.5 | 51.8 | 49.4 | ||||
| (b) | 51.7 | 62.1 | 52.5 | 50.3 | ||||
| (c) | 56.9 | 68.3 | 57.6 | 55.4 | ||||
| (d) | 58.7 | 69.2 | 58.5 | 56.8 | ||||
| (e) | 61.8 | 71.9 | 60.8 | 58.3 | ||||
IV-C Ablation Studies
IV-C1 The number of global and attribute-wise representations
Global and attribute-wise representations capture individual nuances in intra-domain and inter-domain consistency. We evaluate the suitability of multiple global and attribute-wise representations as shown in TABLE V. We have observed that setting the number of global and attribute-wise representations to 3 achieves the best performance for our method.
IV-C2 Performance of Different Components
To evaluate the contribution of each component to our DCR model, we conduct ablation studies on both seen and unseen domains, as detailed in Table VI. Here, PFM denotes the Parallel Fusion Module, while ACN refers to the Attribute Compensation Network. AF and KC represent the attribute-oriented anti-forgetting and knowledge consolidation strategies, respectively. Comparing (a) and (b), we observe that PFM improves performance by +1.3% mAP / +0.6% R-1 on seenavg and +0.7% mAP / +0.9% R-1 on unseenavg under the baseline method. This demonstrates its effectiveness in fusing text and image information to generate diverse global representations. Comparing (a) and (c), integrating PFM and ACN yields significant gains: +6.5% mAP / +6.8% R-1 on seenavg and +5.8% mAP / +6.0% R-1 on unseenavg, for generating the diverse global representations. Comparing (c) and (e), when integrating AF and KC, the model achieves gains of +4.9% mAP / +3.6% R-1 on seenavg and +3.2% mAP / +2.9% R-1 on unseenavg. These strategies significantly promote inter-domain consistency and knowledge transfer. Our DRE method unifies PFM, ACN, AF, and KC into an end-to-end LReID model, striking a balance between maximizing intra-domain discrimination and minimizing inter-domain gaps. These combinations ensure strong generalization while mitigating catastrophic forgetting in our DCR.
@@ IV-C2 Performance of Different Components
| Confidence Threshold | SeenAvg | UnseenAvg | ||
| mAP | R-1 | mAP | R-1 | |
| 0.7 | 60.9 | 71.1 | 59.7 | 57.6 |
| 0.8 | 61.8 | 71.9 | 60.8 | 58.3 |
| 0.9 | 62.0 | 71.6 | 60.5 | 58.0 |
IV-C3 Performance of attribute-text generator
@@
To better understand whether each instance’s text descriptions generated by the attribute-text generator (ATG) provide more fine-grained guidance for learning global representations, we train our model using the generic text descriptor ”A photo of a person” (w/o ATG) for comparison. TABLE VII shows that the attribute-text generator obtains text descriptions to significantly improve overall performance. When using the specific text descriptors, the average decreases by 1.7% mAP/1.4% R-1 on the seen domain and by 1.5% mAP/1.8% R-1 on the unseen domain. ATG enhances the robustness of global representations for each instance, effectively mitigating the forgetting of old knowledge.
IV-C4 Analysis of confidence threshold
To evaluate the sensitivity of the confidence threshold in an attribute-text generation, we conduct an ablation study with threshold values of 0.7, 0.8, and 0.9, as shown in TABLE VIII. The results demonstrate that increasing the confidence threshold from 0.7 to 0.8 consistently enhances DCR performance across metrics. While a confidence threshold of 0.9 achieves the highest 62.0% mAP on seenavg, other metrics do not reach their peak performance at this value. Overall, we find that setting the confidence threshold to 0.8 strikes the best balance, ensuring robust classification accuracy while maximizing model performance.
IV-D Visualization
To further validate the effectiveness of our DCR, we conduct a qualitative comparison with the DKP method, as shown in Fig. 7. The left column displays challenging query images, including cases with occlusion, blur, side views, and back views. In Fig. 7, correctly matched pedestrians are marked with green boxes, while incorrect matches are highlighted in red. Under occlusion conditions (first and second rows), our DCR accurately retrieves eight out of ten individuals with the same identity, despite the limited visual information available. For blurry queries (third row), the DKP method only correctly matches three pedestrians, whereas our DCR achieves perfect retrieval (10/10) even with the unclear query image. Additionally, in side-view and back-view scenarios (fourth and fifth rows), DCR successfully handles clothing color variations, outperforming DKP in matching the same identity. These results demonstrate that DCR achieves superior retrieval performance, consistently identifying more pedestrians with the same identity across diverse challenging conditions. This advantage stems from the collaborative design of global and attribute-wise representations, which balances intra-domain discrimination and inter-domain gaps. By enhancing discrimination between similar identities and adapting to complex scenarios, our DCR demonstrates significantly more robustness than existing approaches.
IV-E Analysis of Speed
We evaluate the computational efficiency and performance of various methods on the Market1501 dataset, which consists of a query set (750 IDs, 3,368 images) and a gallery set (751 IDs, 15,913 images). All experiments were conducted using two NVIDIA RTX A6000 GPUs for training and a single GPU for inference. As shown in Table IX, where time measurements are indicated in hours (h), minutes (min), and seconds (s), our DCR completes training in 19 hours and 10 minutes while achieving an inference time of 65 seconds, processing gallery images at a rate of 52 images per second. Although our DCR shows longer training and inference times compared to AKA, PatchKD, and DKP approaches, it delivers significantly superior re-identification performance, achieving an seenavg incremental gain of 10.0% mAP/7.8% R-1, and 9.8% mAP/7.5% R-1, an seenavg improvement 10.0% mAP/7.8% R-1, and 9.8% mAP/7.5% R-1 on training order-1 and order-2, as demonstrated in Tables III and IV. The increased computational complexity of DCR stems from its transformer-based CLIP architecture incorporating multiple class tokens and the additional attribute compensation network (ACN) that enhances discrimination of similar identities through local feature integration. While this design choice results in higher computational demands, it achieves an excellent balance between performance metrics and complexity. For future work, we plan to explore prompt tuning [53] and mixture-of-experts (MOE) approaches [55] to optimize the model’s efficiency while maintaining its competitive performance advantages.
V Conclusions
@@
In this paper, we propose a domain consistency representation learning (DCR) model that explores global and attribute-wise representations to capture subtle nuances in intra-domain and inter-domain consistency, achieving a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps. Specifically, global and attribute-wise representations serve as complementary information to distinguish similar identities within the domain. We also develop an attribute-oriented anti-forgetting (AF) strategy and a knowledge consolidation (KC) strategy to minimize inter-domain gaps and facilitate knowledge transfer, enhancing generalization capabilities. Extensive experiments demonstrate that our method outperforms state-of-the-art LReID methods.